Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 219     Entry time: 05 Dec 2024, 19:47
Author: Alexey Zozulya 
Group: MID 
Subject: Stability Issue 

Today we have observed multiple large disturbances to beam stability. These can be characterised by large jumps in the horizontal beam position, which at its extreme amounts to a loss of beam from our behalf. M2 mirror feedbacks are able to return the beam to the initial position after ca. 10 minutes. Beam jumps have been approximately 1hr apart. We do not observe the beam on the XTD6 tunnel imager in this case, but do observe good lasing quality on the XGM.

Attachment 1: 2024-12-05_19-43-31.png  472 kB  | Hide | Hide all
2024-12-05_19-43-31.png
ELOG V3.1.4-7c3fd00